given the linear equation y=-3x-2 find the y- coordinates of the points (0)(-3)(-1) please show all of your work. plot these points and graph the liner equation
Answer by jim_thompson5910(35256)

For (0, ?), x = 0. So y=-3x-2 becomes y = -3(0)-2 = -2

So when x = 0, y = -2 and the point is (0, -2)

Do the same for the rest and plot them to graph this line.
-------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

-------------------------------------------------------------------------------------------------